OpenStudy (anonymous):

If a chord of circle with radius 5 cm subtends a right angle at the centre , then length of chord is

OpenStudy (anonymous):



OpenStudy (anonymous):

use Pythagorean theorem then..........

OpenStudy (anonymous):

5^2 + 5^2 = c^2 25+25 = c^2 50 = c^2 c = sqrt(50)

OpenStudy (anonymous):

Yes the answer would be radical 50. Simplify radical(2-25) =\[5\sqrt{2}\]